OpenStudy (anonymous):

can someone please explain to me how you do something like(a-1)(a+1)(n^2+1)? I'm only in year 7 so I've only done (a-1)(1+1) and i know thats called quadratics but when there's 3 brackets, what is it called? Thanks a lot :D

OpenStudy (anonymous):

\[(a-1)(a+1)(n^2+1)\] or \[(n-1)(n+1)(n^2+1)\]??

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

just make it n

OpenStudy (anonymous):

\[(n−1)(n+1)(n^2+1)\] We know: \[(n-1)(n+1)=(n^2-1^2)=(n^2-1)\] \[(n^2-1)(n^2+1)=(n^4-1^2)=(n^4-1)\]

OpenStudy (anonymous):

shouldnt (n-1)(n+1) be n^2-1?

OpenStudy (anonymous):

It is.

OpenStudy (anonymous):

but u said (n-1)(n+1)=(n^2+1^2)

OpenStudy (anonymous):

I said (n^2-1^2) and 1^2=1*1=1